OpenStudy (anonymous):

Find the probability of getting two heads and one tail on a toss of three fair coins

OpenStudy (cwrw238):

possible outcomes are HHT, HTH, THH probability of H or tail n one toss is 1/2 probability of each of the above 3 outcomes = 1/2 * 1/2 * 1/2 probability of one of the possible out are now added

OpenStudy (cwrw238):

last line should read probability of one of the possible outcomes are now added

OpenStudy (cwrw238):

do you follow that?

OpenStudy (cwrw238):

when events are independent the probabilities are multiplied: eg i throw H AND T AND H probability of this is 1/2 * 1/2 * 1/2 AND = multiply where events are mutually exclusive probabilities are ADDED HTH OR THH = 1/8 + 1/8 OR = ADD

OpenStudy (cwrw238):

you still there?

OpenStudy (anonymous):

Sorry, yes I understood. Thank you for your help :)

OpenStudy (cwrw238):

yw